# Beyond Text: Improving LLM’s Decision Making for Robot Navigation via Vocal Cues

### v1

@@ 6 Experimental Results
To validate the efficacy of our approach, we utilize confidence score and winning rate to show that Beyond Text can achieve low variance and low bias results compared to current LLM-based human instructions without vocal cues.
@@ 6.1 Confidence Score Improvement
As discussed in Section 3, we show that our method achieves a low variance with an increased confidence score. To better visualize the low variance of our approach, we first selected a subset of 100 data instances wherein the choices made by Large Language Models (LLMs) were in alignment with user selections, shown in Figure 3. Then, we analyze the average confidence score across all 500 data clips in the dataset to demonstrate the efficacy and generalizability of our framework, shown in Figure 4. Both Figure 3 and Figure 4 revealed a marked increase in confidence score: Beyond Text significantly surpasses text-only language model, both the transcription-only and the with reasoning LLM. This trend was also observed across two distinct categories. The introduction of affective cues significantly lowers variance, enhancing existing LLMs’ ability to interpret human vocal uncertainty with confidence.
Figure 3: Average Confidence Score by context and audio type for samples that language model picks the same choices as human perception.
Figure 4: Average Confidence Score by context and audio type. An overall low variance is indicated by the increased average confidence.
@@ 6.3 Winning Rate
Transcription Only 22.16% 22.79% 21.75%
With Reasoning 49.30% 36.74% 58.60%
@@
6.4 Vocal Cue Ablation Study
To elucidate the impact of individual components within the vocal cue framework, we conduct a comprehensive qualitative winning rate study by isolating each category of vocal cues and presenting them exclusively to the language model. We report the vocal cue ablative study in Table 2.
In the absence of reasoning, the highest winning rate was observed with the exclusive use of loudness features. This suggests that incorporating multiple vocal features without reasoning may overwhelm the model, hindering its ability to correlate vocal characteristics with uncertainty. Conversely, with reasoning integrated, vocal cues significantly strengthened the LLM’s proficiency in interpreting vocal uncertainty and facilitating robotic navigation, as reflected in the increased winning rates.
@@ 7 Robustness to Adversarial LLM Attacks
We demonstrate that Beyond Text exhibits robustness against token manipulation adversarial attacks targeting LLMs. Our attack pipeline is designed as follows: Post-transcription, a paraphraser alters the original transcription ($T1$) into a new version ($T2$), systematically replacing uncertainty semantics with deterministic language. We then adjust the prompt to align the vocal and textual information with $T2$. Subsequently, the LLM is instructed to replicate $T2$ as option (A) instead of generating a freeform paraphrase ($T1^{\prime}$) devoid of uncertainty cues. This approach highlights the limitations of current LLMs, which tend to over-rely on semantic textual instructions with no clue about the nuances of vocal delivery. Figure 6 illustrates the attack pipeline.
@@ 8 Conclusions and Open Questions
(1) Improving Quality of Audio Augmentation: It is essential to consider more audio features, such as intonation and prosody, with learning-based affective computing methods to improve the quality of audio augmentation. The DNIA is a foundational dataset designed for studying human uncertainty in audio-guided navigation tasks, but it can also be used in other HRI-related tasks. In the future, the dataset will be expanded through the addition of more related vocal features. Additionally, the relation between different vocal features also needs to be carefully articulated for different tasks so that LLMs can provide concrete reasoning.
(2) Fine-Grained Definition of Uncertainty: Beyond Text marks a notable first step in interpreting human uncertainty in human-robot conversations, but it has limitations. For instance, a directive like “Turn left at the washroom. Oh, sorry, turn right at the washroom” could signify either uncertainty or a corrected command, highlighting the complexity of human speech. It is crucial to address such uncertainty in a fundamental manner in future research.
(3) Advancing LLMs to Listen Audio without Text: While we currently rely on textual transcriptions, emerging research, notably (Tang et al., 2023), demonstrates LLMs processing audio directly with considerable success in speech tasks. However, these models still lack the capability to fully grasp human emotional nuances, especially in conveying uncertainty.
@@
9 Impact Statements
This paper aims to contribute to the development of robotics systems powered by language models, offering preliminary findings that enhance their reliability and confidence. These initial outcomes lay the groundwork for establishing dependable and robust language-based robotics systems capable of executing tasks effectively while listening to human feedback, thus benefiting society at large. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A LLM Implementation Design
We present the pipeline of querying LLM to analyze human audio inputs. GPT-4 is used for choice generation in the first stage with in-context examples provided. The response contains the chain-of-thought reasoning process for the final action choice. The LLM output is cascaded to GPT-3.5 for scoring the confidence of each choice. In each stage, we enforce LLM to return output in JSON format, ensuring consistency in both our implementation and the parsing process. We also encapsulate the prompts in an API request parallel processor to speed up our workflow.
@@
Appendix B Example Prompts and LLM Outputs
B.1 Prompts and Examples for Generating Reasoning and Choices (GPT-4)
You are a robot exploring a building. You are in the middle of the building and are seeking directions from a human. Your goal is to provide possible next steps based on human response. Human response may not be trustworthy and contain disfluency. You should figure out the unreliable part of the human response and generate five different next-step actions based on the reasoning. Your answer must strictly follow the JSON format the example below provided. You must strictly follow the format of the example provided below for your reasoning and next-step actions. Each next-step action must be distinct. The last next-step action, E), must be ”Ask another person near you for direction.” For the rest next-step actions, if there is uncertainty, say ”ask other people for further instruction” at the instruction where the uncertainty appears. Do not paraphrase. Always provide reasoning.
@@ B.1.1 Without Affective Cue
    "Reasoning":"The use of maybe indicates uncertainty in the human response, which makes it not fully trustworthy.",
    "Answer":{
    A: Start from the mail office, walk straight and take the first right turn, then go straight and you will see the cafeteria
    B: Start from the mail office, walk straight and take the second right turn, then go straight and you will see the cafeteria
    C: Start from the mail office, walk straight and take the first left turn, then go straight and you will see the cafeteria
    D: Start from the mail office, walk straight and at the first right turn, ask another person for further instruction
    E: Ask another person near you for direction
B.1.2 With Affective Cue
@@
    "Reasoning":"The phrase "And then you will probably need to make a left turn" has a long duration of 7.40 seconds, which indicates hesitation or uncertainty in the direction given. Additionally, the large pitch change at 23.000 second, which falls within the duration of "and you will see lab B", suggests that the speaker is unsure about the final part of the instruction. The use of the word probably also adds to the uncertainty of the direction to make a left turn.",
    "Answer":{
    A: Start from lab A, go straight, pass lab A and the locker room, make a left turn, go straight, and you will see lab B
    B: Start from lab A, go straight, pass lab A and the locker room, and ask other people for further instruction
    C: Start from lab A, go straight, pass lab A and the locker room, make a right turn, go straight, and you will see lab B
    D: Start from lab A, go straight, pass lab A and the locker room, make a left turn, and then ask other people for further instruction
    E: Ask another person near you for direction
B.2 Prompts and Examples for Choice Selection (GPT-3.5)
@@ B.2.1 Transcript Only
B.2.2 With Transcription and Reasoning
@@
B.2.3 Transcription, Reasoning, and Affective Cue
@@
B.3 Large Language Model Adversarial Token Manipulation Attack Prompt
The LLM attack use GPT4 as the adversarial paraphraser to attack the original text input. The attack takes the sub-instruction list of each original audio clip as input, then paraphrase it to reverse the semantic tone. The GPT4 attack prompt is shown below:
@@ B.3 Large Language Model Adversarial Token Manipulation Attack Prompt
Appendix C Transcription Model Prompting
We employ the state-of-art Whisper model (Radford et al., 2023) to transcribe the human audio-guided instruction. We use Whisper-Small with 244M parameters that required  2GB VRAM to maintain the seamless transcription speed (within 1 second). We add a custom prompting ”Umm, let me think like, hmm… Okay, here’s what I’m, like, thinking.” to enhance its proficiency in capturing subtle nuances of speech, such as word mumbling, hesitations, and speech repairs that may sounds like a slips of tongue, as these are crucial uncertainty-related semantic signals.
Appendix D Pitch and Loudness Shifts Detection
Algorithm 1 detects largest shifts in pitch and loudness in an given audio clip. We use the timestamp to align with the sequence of instructions for vocal cue prompts.
@@ Appendix D Pitch and Loudness Shifts Detection
0:  audio, $\theta_{l}$ (loudness threshold), $\theta_{p}$ (pitch threshold)
@@
  for each second $t_{i}$ in audio do
@@
Appendix E User Interface for Human Evaluation on DNIA Dataset
This section details the user interface that we used for human assessment. As shown in Figure 8, the website allows user to login and participate the human evaluation session. The user is asked to listen to an randomly selected audio clip from the audio dataset and answer the question by choosing the most accurate and efficient answer from a list of five options. Once the user select the answer, they need to hit Confirm button to submit the answer the proceed to the next question.

### v2

@@ 5 Experimental Results
To validate the efficacy of our approach, we utilize confidence score and winning rate to show that Beyond Text can achieve low variance and low bias results compared to current LLM-based human instructions without vocal cues. We will first introduce the evaluation metric, and the comparison experiments, ablation studies on various vocal cues and LLM types.
5.1 Evaluation Metric: Confidence Measure
$\mathcal{C}(\rho)=\frac{1}{\text{KL}(\rho,\rho^{*})},$ (3)
where we note that the confidence measure will reduce as the KL divergence between $\rho$ and $\rho^{*}$ increases. The range of a confidence measure, which is inversely proportional to entropy, is bounded between $[0,\infty]$. Ideally, we would like to have the confidence measure as high as possible, which requires us to reduce $\text{KL}(\rho,\rho^{*})$. We first expand and note that
where $H(\rho)=-\sum_{j=1}^{J}\rho(y^{j})\log(\rho{(y^{j})})$ is the entropy of predicted distribution $\rho(y^{j})$ by the LLM which quantifies the amount of information contained in random variables. We note that since we don’t have any control over the second term in equation 4 because $\rho^{*}$ is unknown, we can focus on minimizing the first term in equation 4, which tells us to maximize the entropy in order to reduce the KL divergence between $\rho$ and $\rho^{*}$. We hypothesize that one of the major reasons for this uncertainty is the information loss while considering only textual information $x$ and losing various critical attributes of the input audio information $v$. From the standard information theory, one can quantify this loss in information as $\mathcal{I}_{l}=H(V)-I(X;V)$ as the difference between the entropy of the original audio input random variable $V$ and the mutual information between the text random variable $X$ and $V$ represented as $I(Y;V)$ which measures the amount of information that is shared between $X,V$. Notably, a higher confidence score—which indicates a lower KL divergence between $\rho$ and $\rho^{*}$—is desirable, as it signifies a closer alignment with the ground truth. In this work, we empirically show that adding the vocal cue in the input helps to reduce the $\text{KL}(\rho,\rho^{*})$ by plotting the confidence measure equation 3 for robotic navigation tasks in Section 5.
@@ 5.2 Confidence Score Improvement
As discussed in Section 5.1, we show that our method achieves a low variance with an increased confidence score. To better visualize the low variance of our approach, we first selected a subset of 100 data instances wherein the choices made by Large Language Models (LLMs) were in alignment with user selections, shown in Figure 3. Then, we analyze the average confidence score across all 500 data clips in the dataset to demonstrate the efficacy and generalizability of our framework, shown in Figure 3. It reveals a marked increase in confidence score: Beyond Text significantly surpasses the text-only language model, both the transcription-only and the LLM using Chain-of-Thought (CoT). This trend was also observed across two distinct categories. The introduction of affective cues significantly lowers variance, enhancing existing LLMs’ ability to interpret human vocal uncertainty with confidence.
@@ 5.4 Winning Rate
Transcription Only LLM 22.16% 22.79% 21.75%
With Chain-of-Thought Reasoning 49.30% 36.74% 58.60%
@@
Table 4: Winning Rate on Different LLMs
LLMs Winning Rate
All VU LU
GPT-3.5-turbo-instruct (Transcription Only) 22.16% 22.79% 21.75%
GPT-3.5-turbo-instruct (Beyond Text) 70.46% 72.56% 68.77%
text-davinci-003 (Transcription Only) 57.88% 57.67% 57.89%
text-davinci-003 (Beyond Text) 69.06% 69.77% 68.42%
Gemini-1.5-Pro (Transcription Only) 43.11% 41.40% 44.56%
Gemini-1.5-Pro (Beyond Text) 65.27% 66.98% 63.86%
5.5 Ablation Study
Vocal Cue: To elucidate the impact of individual components within the vocal cue framework, we conduct a comprehensive qualitative winning rate study by isolating each category of vocal cues and presenting them exclusively to the language model. We report the vocal cue ablative study in Table 2. In the absence of reasoning, the highest winning rate was observed with the exclusive use of loudness features. This suggests that incorporating multiple vocal features without reasoning may overwhelm the model, hindering its ability to correlate vocal characteristics with uncertainty. Conversely, with reasoning integrated, vocal cues significantly strengthened the LLM’s proficiency in interpreting vocal uncertainty and facilitating robotic navigation, as reflected in the increased winning rates.
LLM Type: To show the generalized ability of Beyond Text, we test it on three different popular LLMs, including GPT3.5, GPT 3, and Gemini. Based on the result shown in Table 4, we conclude that Beyond Text improves the winning rate on all three LLMs significantly. Details about parameters for each type of LLM is shown in Appendix A.
@@ 6 Robustness to Adversarial LLM Attacks
We demonstrate that Beyond Text exhibits robustness against token manipulation adversarial attacks targeting LLMs. Our attack pipeline is designed as follows: Post-transcription, a paraphraser alters the original transcription ($T1$) into a new version ($T2$), systematically replacing uncertainty semantics with deterministic language. We then adjust the prompt to align the vocal and textual information with $T2$. Subsequently, the LLM is instructed to replicate $T2$ as an option (A) instead of generating a freeform paraphrase ($T1^{\prime}$) devoid of uncertainty cues. This approach highlights the limitations of current LLMs, which tend to over-rely on semantic textual instructions with no clue about the nuances of vocal delivery. Figure 5 illustrates the attack pipeline A detail example is shown in Appendix D.
@@ 7 Conclusions and Open Questions
(1) Improving Quality of Audio Augmentation: It is essential to consider more audio features, such as intonation and prosody, with learning-based affective computing methods to improve the quality of audio augmentation. The DNIA is a foundational dataset designed for studying human uncertainty in audio-guided navigation tasks, but it can also be used in other HRI-related tasks. In the future, the dataset will be expanded through the addition of more related vocal features. Additionally, the relation between different vocal features also needs to be carefully articulated for different tasks so that LLMs can provide concrete reasoning. As we stated the position that affective cue analysis is currently missing from existing LLM, and beyond text is an audio augmentation method that can help LLM hear not only ‘what’ people said but also ‘how’ people convey their ideas; it is also excited to explore the potential of vocal features in improving audio-to-text tasks, such as video emotion understanding, etc.
(2) Fine-Grained Definition of Uncertainty: Beyond Text marks a notable first step in interpreting human uncertainty in human-robot conversations, but it has limitations. For instance, a directive like “Turn left at the washroom. Oh, sorry, turn right at the washroom" could signify either uncertainty or a corrected command, highlighting the complexity of human speech. It is crucial to address such uncertainty in a fundamental manner in future research.
(3) Advancing LLMs to Listen Audio without Text: While we currently rely on textual transcriptions, emerging research, notably Tang et al. (2023), demonstrates LLMs processing audio directly with considerable success in speech tasks. However, these models still lack the capability to fully grasp human emotional nuances, especially in conveying uncertainty.
@@ Appendix A LLM Implementation Design
We present the pipeline of querying LLM to analyze human audio inputs in Figure 6. GPT-4 is used for choice generation in the first stage, with in-context examples provided. The response contains the chain-of-thought reasoning process for the final action choice. The LLM output is cascaded to GPT-3.5 for scoring the confidence of each choice. In each stage, we enforce LLM to return output in JSON format, ensuring consistency in both our implementation and the parsing process. We also encapsulate the prompts in an API request parallel processor to speed up our workflow.
@@
For clarity and transparency, we used gpt-4-1106-preview for choice generation and gpt-3.5-turbo-instruct model. For Gemini-1.5-Pro and gpt-4-1106-preview, we used default hyperparameters except for the temperature, which is set to 0. For gpt-3.5-turbo-instruct, text-davinci-003, we set max_token: 100; temperature: 0; echo: true; and the following hyper-parameters:
Table 6: Logit Bias Settings
Token ID Character (with space) Logit Bias
362 A 100.0
426 B 100.0
356 C 100.0
423 D 100.0
469 E 100.0
Appendix B Transcription Model Prompting
We employ the state-of-the-art Whisper model (Radford et al., 2023) to transcribe the human audio-guided instruction. We use Whisper-Small with 244M parameters that require 2GB VRAM to maintain the seamless transcription speed (within 1 second). We add a custom prompting "Umm, let me think like, hmm… Okay, here’s what I’m, like, thinking." to enhance its proficiency in capturing subtle nuances of speech, such as word mumbling, hesitations, and speech repairs that may sound like a slip of the tongue, as these are crucial uncertainty-related semantic signals.
Appendix C Example Prompts and LLM Outputs
C.1 Prompts and Examples for Generating Reasoning and Choices (GPT-4)
You are a robot exploring a building. You are in the middle of the building and are seeking directions from a human. Your goal is to provide possible next steps based on human response. A human response may not be trustworthy and contain disfluency. You should figure out the unreliable part of the human response and generate five different next-step actions based on the reasoning. Your answer must strictly follow the JSON format the example below provided. You must strictly follow the format of the example provided below for your reasoning and next-step actions. Each next-step action must be distinct. The last next-step action, E), must be "Ask another person near you for direction." For the rest of the next actions, if there is uncertainty, say "ask other people for further instruction" at the instruction where the uncertainty appears. Do not paraphrase. Always provide reasoning.
@@ C.1.1 Without Affective Cue
{ "Reasoning":"The use of maybe indicates uncertainty in the human response, which makes it not fully trustworthy.", "Answer":{ A: Start from the mail office, walk straight and take the first right turn, then go straight and you will see the cafeteria B: Start from the mail office, walk straight and take the second right turn, then go straight and you will see the cafeteria C: Start from the mail office, walk straight and take the first left turn, then go straight and you will see the cafeteria D: Start from the mail office, walk straight and at the first right turn, ask another person for further instruction E: Ask another person near you for direction } }
C.1.2 With Affective Cue
@@
{ "Reasoning":"The phrase "And then you will probably need to make a left turn" has a long duration of 7.40 seconds, which indicates hesitation or uncertainty in the direction given. Additionally, the large pitch change at 23.000 second, which falls within the duration of "and you will see lab B", suggests that the speaker is unsure about the final part of the instruction. The use of the word probably also adds to the uncertainty of the direction to make a left turn.", "Answer":{ A: Start from lab A, go straight, pass lab A and the locker room, make a left turn, go straight, and you will see lab B B: Start from lab A, go straight, pass lab A and the locker room, and ask other people for further instruction C: Start from lab A, go straight, pass lab A and the locker room, make a right turn, go straight, and you will see lab B D: Start from lab A, go straight, pass lab A and the locker room, make a left turn, and then ask other people for further instruction E: Ask another person near you for direction } }
C.2 Prompts and Examples for Choice Selection (GPT-3.5)
@@ C.2.1 Transcript Only
C.2.2 With Transcription and Reasoning
@@
C.2.3 Transcription, Reasoning, and Affective Cue
@@
C.3 Large Language Model Adversarial Token Manipulation Attack Prompt
The LLM attack uses GPT4 as the adversarial paraphraser to attack the original text input. The attack takes the sub-instruction list of each original audio clip as input and then paraphrases it to reverse the semantic tone. The GPT4 attack prompt is shown below:
@@ C.3 Large Language Model Adversarial Token Manipulation Attack Prompt
Appendix D Adversarial Attack Example
Before Attack:
After Attack:
"First, start from lab A, you go straight and pass lab A and then the locker room. And then you will make a left turn. And then you go straight and you will see lab B." Large loudness decrease: No Change Large pitch change: at time = 23.000 second Duration: "First," => [0.000, 2.160] (2.16 seconds); Duration: "start from lab A," => [2.160, 4.200] (2.04 seconds); Duration: "you go straight" => [4.200, 6.840] (2.64 seconds); Duration: "and pass lab A" => [6.840, 9.280] (2.44 seconds); Duration: "and then the locker room." => [9.280, 12.440] (3.16 seconds); Duration: "And then you will make a left turn." => [12.440, 19.840] (7.40 seconds); Duration: "And then you go straight" => [19.840, 22.720] (2.88 seconds); Duration: "and you will see lab B." => [22.720, 28.080] (5.36 seconds););
Appendix E DNIA Dataset Details and Samples
For VU clips, we ensure no semantic uncertainty signals, such as uncertain-related words (i.e., ‘maybe,’ ‘I guess’), speech repairs, and hesitations. We believe those are common affective characteristics associated with human speech. However, those affective cues are not prominent in LU clips. As our position states that affective cues (duration, pitch, loudness) are associated with speech uncertainty levels in robot navigation tasks, we believe the setup of our dataset is comprehensive enough to show that vocal cues are good predictors. For review purposes, we create an anonymous GitHub repo: https://anonymous.4open.science/r/DNIA-dataset-D22E. The full dataset will be publicly released soon after the review as we prepare them for storage and access control. The dataset will come with textual annotations (grounded labels) as well as Whisper-generated transcription to save resources for other researchers who can reproduce the result.
Appendix F Pitch and Loudness Shifts Detection
Algorithm 1 detects the largest shifts in pitch and loudness in a given audio clip. We use the timestamp to align with the sequence of instructions for vocal cue prompts.
@@ Appendix F Pitch and Loudness Shifts Detection
0:  audio $\mathcal{V}$, $\theta_{l}$ (loudness threshold), $\theta_{p}$ (pitch threshold)
@@
  for each second $t_{i}$ in $\mathcal{V}$ do
@@
Appendix G User Interface for Human Evaluation on DNIA Dataset
This section details the user interface that we used for human assessment. As shown in Figure 7, the website allows users to log in and participate in the human evaluation session. The user is asked to listen to a randomly selected audio clip from the audio dataset and answer the question by choosing the most accurate and efficient answer from a list of five options. Once the user select the answer, they need to hit Confirm button to submit the answer the proceed to the next question.